OpenStudy (anonymous):

Please help!!!!! MEDAL AND FAN!!!!! Give the center coordinates (x,y) of the circle:

OpenStudy (anonymous):

\[(x+3)^{2}+(y-4)^{2}=36\]

OpenStudy (triciaal):

equation of a circle (x-h)^2 + (y-k)^2 = r^2 center (h,k) and radius r

OpenStudy (anonymous):

Center = (-3,4) This is because, as triciaal said, in the equation (x+3)2+(y−4)2=36, 3=h and 4=k (H,K) are the coordinates of the center.

OpenStudy (anonymous):

Oh ok

OpenStudy (amilapsn):

The definition of the circle is: the loci of a point which is moving at constant distance from a fixed point. So if the point is (h, k) and the constant distance is r, by the distance formula: \[\sqrt{(x-h)^2+(y-k)^2}=r\] After squaring, equation of the circle: \[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

When multiplied out, we obtain the "general form" of the equation of a circle. Notice that in this form we can clearly see that the equation of a circle has both x2 and y2 terms and these terms have the same coefficient (usually 1).

OpenStudy (anonymous):

Wait so which explanation should I go with because I a bit confused

OpenStudy (anonymous):

@ikram002p

OpenStudy (triciaal):

all are correct! different amount of details.

OpenStudy (anonymous):

ok thx

OpenStudy (anonymous):

yeah there all the same

OpenStudy (anonymous):

But..... the first time I did this problem, I put -3,4 for my answer and got it wrong

OpenStudy (freckles):

ordered pairs should be in the form (h,k) that is you should have (-3,4)

OpenStudy (freckles):

Maybe they already included the ( ). Don't know.

OpenStudy (anonymous):

oh ok I will put it in parentheses

OpenStudy (anonymous):

Could y'all help me on on more problem?

OpenStudy (anonymous):

*one

OpenStudy (anonymous):

Sure!

OpenStudy (anonymous):

\[(x + 1)^{2} + (y - 6)^{2}= 2.25\]

OpenStudy (anonymous):

Would the coords be (-1,6)?

OpenStudy (freckles):

yes the center would be (-1,6)

OpenStudy (anonymous):

Absolutly right!

OpenStudy (anonymous):

thx